OpenStudy (anonymous):

Determine the improper fraction that is equivalent to the mixed number. 28\tfrac{3}{5} A. \frac{84}{5} B. \frac{143}{5} C. \frac{90}{5} D. \frac{31}{5}

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

28 3/5=(28*5+3)/5 simplify this and get your answer

OpenStudy (anonymous):

I'm horrible at simplyifing and I have no idea what you just wrote.

OpenStudy (alexandervonhumboldt2):

\[28\frac{ 3 }{ 5 }=\frac{ 28*5 +3 }{ 5 }\]

OpenStudy (alexandervonhumboldt2):

simplify the last fraction

OpenStudy (alexandervonhumboldt2):

it equals to 143/5

OpenStudy (anonymous):

Wait. So I am simplyfying 143/5? how?

OpenStudy (alexandervonhumboldt2):

no...... it is already simplified

OpenStudy (anonymous):

o